how can u change the equation "x - 5 y = 103" into a "y = mx + b" format.
Answer by nerdybill(7384)

If you started with:
x - 5 y = 103
.
Start by moving the 'x' to the right. Do this by subtracting an 'x' from each
side.
x - 5 y - x = -x + 103
-5y = -x + 103
.
Now, divide both sides by -5:
(-5y)/(-5) = [-x + 103]/(-5)
y = (1/5)x + (103/-5)
y = (1/5)x - (103/5)
That's it, you're done.
.
slope (m) is (1/5)
y-intercept (b) is at (0, -103/5)